# Dynamics of Complex Quadratic Families under Holomorphic and Nonholomorphic Singular Perturbations

### v2

@@ 1 Introduction
It is well known that differential equations can be applied to model many phenomena in nature. Especially after Newton’s fundamental discovery successfully explained the motion of a two-body system, this viewpoint was widely accepted by most scientists. In the following two centuries, Newton’s successors carried on his methodology and tried to construct a unified linear model to interpret the whole natural world. However, all attempts to explicitly and analytically solve the three-body problem ended up with failure. Some physical scientists gradually realized the flaw of the linearization method: the local linearization of nonlinear problems causes the “full truth” of whole systems to be shadowed by the “partial beauty” of local linear properties.
Remark 1.
Because the solutions obey the principle of superposition in linear systems, we can break up a linear equation into smaller (and simpler) pieces and solve them separately, then the combination of these solutions will be a complete solution of the original system. However, for nonlinear systems, we cannot use the same method to find complete solutions.
At the end of 19th century, H. Poincaré suggested an entirely new viewpoint: instead of analytically and explicitly solving the solutions of the three-body system, trying to qualitatively describe the behaviors of the solutions of the differential equations. General speaking, the notions of qualitative behaviour include fixed points (stationary solutions), periodic orbits (periodic oscillations), orbits that show forward or backward asymptotic behaviour, and orbits that show both forward and backward asymptotic behaviour (homoclinic orbits) [1] [2]. H. Poincaré believed that these homoclinic orbits fill up most of the phase space; however, after a few decades this mistake motivated other forms of qualitative analysis, such as recurrent trajectories, instability, mixing, entropy, etc., which collectively influenced the development of chaos.
H. Poincaré’s successor, G. Birkhoff, started to research topological dynamics and ergodic theory [2]. At the same time, J. Julia and P. Fatou began to study the complex dynamics and built the fundamental of this field [3]. In the following thirty years, A. Kolmogorov, V. Arnold and J.Moser well researched the stability in the Hamiltonian systems, which improved G. Birkhoff’s unfinished work and opened a new era in dynamical systems [4]. The progress in complex dynamical systems, however, was relatively slow during this period due to the lack of powerful computation tools [3].
In 1963 E. Lorenz suggested a classical example in his famous article, “Deterministic Non-periodic Flow” [5]. In this paper, he suggested a set of three-dimensional differential equations to describe the convection in the atmosphere, and then found that the simulation result is sensitive to the initial conditions, which was called the “Butterfly Effect” later. In 1971 D. Ruelle and F. Takens proposed a new theory of the turbulence in dissipative systems based on an embryonic consideration of the strange attractors [6]. A few years later, T. Li and J. Yorke published their paper, “Period Three Implies Chaos”[7], which was the first time that the term “chaos” was firstly formally suggested (actually, A. Sarkovskii, claimed a similar thought and gave a more general theorem ahead of Li and Yorke [8]). At the same time, R. May demonstrated the chaotic phenomena hidden behind the changes of population in the Logistic equations [9]. Besides, other physical scientists, such as A. Winfree [10] and M. Feigenbaum [11] [12], also brought more fresh air to this field during this period. This decade is a glorious time of chaos theory. In 1980s, with the rapid development of computer science, more and more theories in dynamical systems (especially complex dynamics) were discovered, and a new branch, “fractal geometry”, was established by B. Mandelbrot [13] based on new numerical analysis and graphic technology.
Remark 2.
Since chaotic motions are non-periodic, we can treat the period of a chaotic system as infinite in some sense. In other words, chaos systems are always related to ``infinity". Therefore, how to tell whether a finite physical process is chaotic or not ((or how to distinguish between a chaotic process and a large-period process)) is a crucial problem. Traditionally, we can say a physical system is chaotic if it is similar enough to the description in the mathematical definition of chaos; in other words, chaotic physical systems not only have positive Lyapunov exponents, but also are sensitive to the initial conditions.
@@
Nowadays, many branches of dynamical systems have been established and well studied, and chaos is treated as one of the most significant scientific discoveries in last century. Although the field of dynamical systems admits a rich variety of viewpoints, the generally accepted objective of dynamical systems involves “the study of asymptotic behaviour of almost all orbits in observation of representative finite parameter families” [1]. In the course of our work, representative simply means those parameter values which produce nontrivial maps. And for the definition of the core concept chaos, a classical one was given by R. Devaney [8], that is: Let $X$ be a metric space and $F:X\rightarrow X$ be a continuous map, then F is chaotic if (1) F is transitive; (2) F is sensitive dependence on the initial conditions; (3) the periodic points of dense in X.
Remark 3.
In these three conditions, “sensitivity” is the central idea in chaos. Actually, “transitivity” and “the density of periodic points” imply sensitivity [14] [15] (see [15] for the proof). Besides, there are some other qualitative or quantitative ways to define chaos, such as topological entropy and power spectrum. See [14] for more discussion. It is worth to mention that ”transitivity” implies the other two conditions in one-dimensional case.
@@
Among different dynamical systems, the dynamics of some types of quadratic families have been well studied and many elegant results have been revealed, while the dynamics of many other types remain unknown and require further research. In this paper, we restrict our discussion to the real and complex quadratic maps. This paper is organized as follows: in the next section, based on some specific examples, we review and summarize the dynamics of some one- and two- dimensional quadratic maps from both topological and statistical viewpoints. In section three, we briefly review some pre-existing results of the dynamics of complex quadratic maps under both holomorphic and nonholomorphic singular perturbations, and the research the dynamics of a special case with an emphasis on the real line. The author’s original research results are mainly presented in section 3.2 and section 3.6. The summary is stated in section four.

### v3

@@ 1 Introduction
It is well known that differential equations can be used to model many phenomena in nature. Following Newton’s fundamental discovery, which successfully explained the motion of a two-body system, this perspective was widely accepted by most scientists. Over the next two centuries, Newton’s successors continued his methodology and attempted to construct a unified linear model to describe the entire natural world. However, all attempts to explicitly and analytically solve the three-body problem ended in failure. Some physical scientists gradually recognized the limitation of linearization: the local linearization of nonlinear problems causes the “full truth” of the system to be overshadowed by the “partial beauty” of local linear properties.
At the end of the 19th century, H. Poincaré proposed an entirely new viewpoint: rather than analytically solving the three-body problem, one should qualitatively describe the behavior of solutions to the differential equations. Generally, qualitative behavior includes fixed points (stationary solutions), periodic orbits (periodic oscillations), orbits exhibiting forward or backward asymptotic behavior, and orbits exhibiting both forward and backward asymptotic behavior (homoclinic orbits) [1, 2]. Poincaré initially believed that these homoclinic orbits occupied most of the phase space; however, this incorrect conjecture eventually inspired other forms of qualitative analysis, such as recurrent trajectories, instability, mixing, and entropy, which collectively influenced the development of chaos theory. This qualitative analysis perspective later evolved into the core paradigm of modern complex systems theory. Complex systems emphasize that the behavior of the whole cannot be deduced solely from the properties of individual components or local interactions; rather, it emerges from the interplay of multiscale interactions and nonlinear couplings within the system. Such systems often exhibit emergent phenomena, where local rules fail to directly account for global patterns. For example, in ecological, economic, or climate systems, the simple behavior of individual components can give rise to highly unpredictable global dynamics. Complex systems theory not only highlights nonlinearity and high dimensionality, but also focuses on adaptability, emergent structures, and self-organization—features that provide an essential conceptual foundation for subsequent developments in chaos theory.
Moreover, the dynamics of complex systems are often constrained by network structures and feedback mechanisms. Positive feedback can amplify local perturbations, leading to cascading effects or sudden transitions, whereas negative feedback tends to stabilize the system or generate oscillatory behavior. Many complex systems also display scale invariance or approximate power-law distributions, implying that the probability of extreme events is far higher than would be expected from exponential decay, and that there exists a nontrivial link between local events and global patterns. By incorporating these concepts, complex systems theory can explain how nonlinear microscopic rules generate diverse dynamical patterns across scales, including chaos, fractals, self-organized criticality, and synchronized or collective behaviors in complex networks. This cross-scale, interdisciplinary framework provides a deeper understanding of modern dynamical systems and chaos theory, and guides the analysis of real-world phenomena that are inherently difficult to predict.
Poincaré’s successor, G. Birkhoff, initiated research in topological dynamics and ergodic theory [2]. Concurrently, J. Julia and P. Fatou began studying complex dynamics and laid the foundations of this field [3]. In the following thirty years, A. Kolmogorov, V. Arnold, and J. Moser conducted in-depth studies on the stability of Hamiltonian systems, building upon Birkhoff’s unfinished work and opening a new era in dynamical systems [4]. Progress in complex dynamical systems during this period was relatively slow due to the lack of powerful computational tools [3]. In 1963, E. Lorenz presented a classical example in his seminal article, “Deterministic Non-periodic Flow” [5]. He proposed a set of three-dimensional differential equations to describe atmospheric convection and discovered that the simulation results were extremely sensitive to initial conditions, a phenomenon later termed the “Butterfly Effect.” In 1971, D. Ruelle and F. Takens proposed a new theory of turbulence in dissipative systems based on preliminary considerations of strange attractors [6]. A few years later, T. Li and J. Yorke published “Period Three Implies Chaos” [7], which marked the first formal introduction of the term “chaos” (although A. Sarkovskii had independently proposed a similar, more general theorem [8]). Simultaneously, R. May demonstrated chaotic phenomena in population dynamics using the logistic equation [9]. Other researchers, such as A. Winfree [10] and M. Feigenbaum [11, 12], also significantly advanced the field during this period. The 1970s thus represented a pivotal decade in the development of chaos theory. In the 1980s, with rapid advances in computer science, many new theories in dynamical systems, particularly complex dynamics, were discovered. B. Mandelbrot established a new branch, “fractal geometry,” based on numerical analysis and graphical technology [13]. Because chaotic motions are non-periodic, the effective period of a chaotic system can be considered infinite. Consequently, chaotic systems are intrinsically linked to the concept of “infinity.” A central problem is determining whether a finite physical process is truly chaotic or simply exhibits very long periodicity. Traditionally, a physical system is considered chaotic if it satisfies the mathematical definitions of chaos, namely, exhibiting positive Lyapunov exponents and sensitivity to initial conditions.
Further studies have revealed that many complex and chaotic systems exhibit power-law distributions, indicating that the probability of extreme events decays much more slowly than in exponential distributions. This means that rare but large-scale events (fat tails) can play a disproportionate role in system behavior. Power-law behaviors imply scale-invariance, where statistical patterns observed at small scales can provide insight into large-scale dynamics. Such scaling properties are closely related to fractal geometry and multi-scale correlations observed in chaotic attractors, making them central to understanding emergent phenomena across physical and biological processes [14, 15], geological history [16, 17, 18], and life evolution [19, 20], and other dynamical systems. Power-law distributions often arise in conjunction with multi-level feedback mechanisms. Positive feedback can amplify small perturbations, triggering cascading effects, tipping points, or abrupt regime shifts, whereas negative feedback constrains fluctuations, maintaining stability or producing oscillatory dynamics. Positive and negative feedback mechanisms widely exist in natural systems and processes [21, 22, 23]. The interplay of these feedbacks can lead to self-organized criticality, a state in which systems naturally evolve toward critical points where minor perturbations produce avalanches of all sizes. Understanding how feedback interacts with nonlinear network dynamics is therefore crucial for linking classical chaos theory with modern complex systems, providing insight into predictability, resilience, and emergent behavior across natural and artificial systems.
@@
Nowadays, dynamical systems theory encompasses a wide range of perspectives, and chaos is recognized as one of the most significant scientific discoveries of the twentieth century. The generally accepted objective involves “the study of asymptotic behaviour of almost all orbits in observation of representative finite parameter families” [1], where “representative” refers to parameter values producing nontrivial maps. R. Devaney [8] provided a classical definition of chaos: let $X$ be a metric space and $F:X\rightarrow X$ a continuous map. Then $F$ is chaotic if (1) $F$ is transitive; (2) $F$ exhibits sensitive dependence on initial conditions; (3) the periodic points are dense in $X$. Among these three conditions, sensitivity is central. In fact, transitivity and the density of periodic points imply sensitivity [24, 25]. Other qualitative or quantitative measures, such as topological entropy and power spectrum, provide additional characterizations of chaos [24]. In one-dimensional systems, transitivity alone implies the other two conditions.
@@
The dynamics of some quadratic families have been extensively studied, revealing many elegant results, whereas the dynamics of other families remain largely unexplored. In this paper, we focus on real and complex quadratic maps. The paper is organized as follows: Section 2 reviews representative one- and two-dimensional quadratic maps from both topological and statistical perspectives. Section 3 summarizes prior results on complex quadratic maps under holomorphic and nonholomorphic singular perturbations and investigates a special case emphasizing the real line, with original contributions presented in Sections 3.2 and 3.6. Section 4 provides a summary.